# Extended Spin-Coherence Time in Strongly-Coupled Spin Baths in Quasi Two-Dimensional Layers

### v1

@@ VI.1 Bath parameters
In our numerical simulations, we first create a very large diamond lattice with the NV-center in the center of this lattice and distribute P1-centers randomly on this lattice. P1-center spins sufficiently far away from the NV spin should have a negligible effect on the spin-coherence decay of the NV spin. Therefore, we limit the number of P1-center spins included in the simulation. We do this by introducing a bath radius $r_{\text{b}}$ which defines a sphere around the NV spin. Only spins within this sphere are considered dynamical.
@@
Figure S4: Evidence for the convergence of parameters for the pCCE(2,4) method applied to a large bath of P1-center electron spins in a quasi two-dimensional configuration at a concentration of 1 ppm: a) the bath-radius $r_{\text{b}}$ and b) the dipole radius $r_{\text{d}}$. The values of $r_{\text{b}}$ and $r_{\text{d}}$ depend both on the layer thickness $L$ and the bath-spin concentration $\rho_{\text{P1}}$. In our calculations, we choose the values for the parameters $r_{\text{b}}$ and $r_{\text{d}}$ corresponding to converged spin-coherence decay results, i.e., the spin-coherence decay does not change appreciably when increasing any of the parameters. For further details, see text.
The values of the above parameters $r_{\text{b}}$ and $r_{\text{d}}$ were determined in convergence studies. An example is illustrated in Fig. S4, where the spin-coherence decay $\langle M_{x}(2\tau)\rangle$ of the NV spin is shown for a P1-center spin bath at concentration 1 ppm using pCCE(2,4). The spin-coherence decay $\langle M_{x}(2\tau)\rangle$ is averaged over 100 randomly distributed spin systems and 20 mean-field spin configurations in internal averaging. The bath radius is increased in steps of 10 nm starting from 60 nm, see Fig. S4 a. The dipole radius is increased in steps of 10% starting from the initial value of 45 nm, see Fig. S4 b. In both calculations, the other parameter was kept constant at its initial value. The resulting graphs do not show any significant changes when varying the parameter values. We thus use the values $r_{\text{b}}=60$ nm and $r_{\text{d}}=45$ nm in our simulations for this particular system. Within the radius $r_{\text{b}}=60$ nm, there are 140 bath spins, which corresponds to the number of P1-centers. Small fluctuations of results shown in Fig. S4 arise presumably due to a finite number of samples for the mean-field average. The chosen parameter values are similar to those used in other studies of NV spins in P1-center baths for CCE2, see for example Ref. [48].
When considering other concentrations $\rho_{\text{P1}}$, the bath radius $r_{\text{b}}$ is chosen such that there are at least 140 spins, i.e., P1-centers, within the radius. To assure partitions of the same size within each P1-bath subgroup (discussed above), further spins are added to each subgroup until the total number of spins is divisible by the partition size $K$.
@@
where the last term on right-hand-side makes $r_{\text{d}}$ larger for quasi two-dimensional layers. We also use the same values of $r_{\text{b}}$ and $r_{\text{d}}$ when leaving out the hyperfine interaction in the P1-centers.
Figure S5: Convergence of the pCCE method with respect to the partition size $K$ and the number of samples for the internal mean-field average for different spin baths. a) An electron spin-1/2 bath in a quasi two-dimensional layer of thickness $L=5$ nm at a concentration of 1 ppm. At larger $K$, results for different numbers of samples for the internal average are also shown to demonstrate the convergence with the number of samples. b) A P1-center spin bath in a quasi two-dimensional layer ($L=30$ nm) at a concentration of $\rho_{\text{P1}}=1$ ppm. Fluctuation of results of pCCE(2,3) likely arise due to insufficient internal mean-field averaging. In a and b, the results for $K=2$ and $K=3$ were averaged internally over 100 and 50 mean-field samples, respectively.
@@ VI.2 Partition size $K$ and the number of samples for the mean-field average
We first consider a quasi two-dimensional system of electron spins-1/2 with layer thickness $L=5$ nm and concentration of 1 ppm and calculate the spin-coherence decay $\langle M_{x}(2\tau)\rangle$. Figure S5 a shows results for different partition sizes $K$ and different number of samples for the internal mean-field average. The results indicate that the convergence is achieved for partition size $K=4$ and 20 samples for the internal mean-field average.
@@
The decay of $\langle M_{x}(2\tau)\rangle$ becomes faster for larger $K$. This implies that, when using CCE2, the resulting $T_{2}$ time is too large in the quasi two-dimensional case. On the other hand, when considering three-dimensional baths of electron spins-1/2, the pCCE(2,$K$) results for all $K\geq 1$ demonstrate similar behaviour (not shown here) which indicates that already pCCE(2,1) (equivalent to CCE2) provides accurate results. For the sake of consistency, we use pCCE(2,4) also for these systems with the same number of samples for the internal average.
@@
For the P1-center spin baths, no significant difference between pCCE(2,1) and pCCE(2,4) was observed for all layers, as indicated in Fig. S5 b. We attribute this to the strong suppression of flip-flop transitions in the bath due to the hyperfine interaction, see the first section of this supplemental material. The CCE2 method thus yields sufficiently accurate results for P1-center spin baths. However, this is strictly speaking valid only for the spin-coherence decay averaged over many different random spin distributions. This does not imply that, for each individual spin system, CCE2 yields sufficiently accurate results, cf. Ref. [53]. Using pCCE, accurate results can also be obtained for individual systems. It is also noteworthy, that, for establishing the validity of CCE2 approch for P1-center spin baths, the extensive use of the pCCE method was necessary.
VII Comparison with the conventional CCE method
We apply CCE3 to the bath of 20 electron spins-1/2 in a quasi two-dimensional layer considered in Fig. 2 of the main article. We vary the number of samples for internal averaging to mitigate the unphysical behavior $\langle M_{x}(2\tau)\rangle>1$; the results are shown in Fig. S6. The results imply that, with increasing number of samples for the internal mean-field average, no enhancement of the accuracy is achieved because the results do not converge even after averaging over 1000 samples. We conclude that this small bath of pure electron spins-1/2 in quasi two dimensions represents a challenging system for the conventional CCE method requiring a large number of samples for the mean-field average.
Figure S6: Results of CCE3 applied to the 20 electron spin-1/2 bath in a quasi two-dimensional layer (from Fig. (2)) for different numbers of samples for the internal average as indicated in the legend. At 1000 samples for the internal average, the CCE3 results have not converged and no signatures of convergence can be observed.
Table S1: Statistics for the occurrence of unphysical behavior ($\langle M_{x}(2\tau)\rangle>1$) for CCE3. We consider a series of smaller systems that we obtain by retaining only the 5, 10 or 15 closest spins to the NV-center from the system of 20 electron spins in a quasi two-dimensional layer, shown in Fig. 2 of the main article. These calculations were repeated 50 times and the fraction indicated in this table refers to this number of repetitions.
Number of spins Occurrence of $\langle M_{x}(2\tau)\rangle$ > 1
5 0 %
10 28 %
15 26 %
20 40 %
We assume that the problem of unphysical behaviour ($\langle M_{x}(2\tau)\rangle>1$) becomes worse with increasing system size because, in larger spin baths, even more clusters are formed, each of which can lead to $\langle M_{x}(2\tau)\rangle>1$. To challenge this assumption, we consider a series of smaller systems that we obtain by retaining only the 5, 10 or 15 closest spins to the NV-center from the system of 20 electron spins considered in Fig. S6. We use CCE3, fix the number of samples for the internal mean-field average to 250 and count the occurrence of unphysical behavior, i.e., $\langle M_{x}(2\tau)\rangle>1$, while repeating the same calculation with random choices of the mean-field bath-spin configurations. The probability for the occurrence of such behavior increases with larger system size as shown in Table S1.
@@ VII Comparison with the conventional CCE method
VIII Fitting the stretched-exponential function
Similar behavior is observed for the baths of P1-centers without the hyperfine interaction, i.e., baths of electron spins-1/2. The resulting curves and corresponding fits are shown in Fig. S8. The transitions from three-dimensional to quasi two-dimensional systems occur at lower layer thickness compared to the P1-center spin bath. The reason for this is the absence of the suppression of the flip-flop transitions by the hyperfine interaction and, hence, different spin dynamics in the bath. Therefore, in Table 1 of the main text, we considered for the quasi two-dimensional configuration for P1-centers a layer thickness of $L=30\,$nm at 2 ppm and 1 ppm and $L=60\,$nm at 0.1 ppm, and, for P1-centers without the hyperfine interaction, $L=5\,$nm at 2 ppm and $L=10\,$nm at 0.1 ppm and 1 ppm.
Figure S8: Illustration of the fitting of the stretched-exponential function in Eq. (2) to the numerical results obtained for P1-centers without the hyperfine interaction for three-dimensional (a) and quasi two-dimensional (b) diamond layers, cf. Table 1 of the main article. The layer thickness $L$ and the P1-center concentration $\rho_{\text{P1}}$ are indicated in the legends. The fitting parameters are the stretched-exponential parameter $p$ and the $T_{2}$ time. The time window for the fit was adjusted for each individual curve separately.
@@ VIII Fitting the stretched-exponential function
The above three regimes complicate the choice for a suitable region for the fit of the short-time decay of $\langle M_{x}(2\tau)\rangle$. Thus, it is difficult to reliably identify the beginning and the end of this region such that the obtained slope of the curves corresponding to the stretched-exponential parameters $p$ together with the characteristic $T_{2}$ time must be considered an approximation. We estimate the error of each individual stretched-exponential parameter $p$ obtained to be at least $\pm$0.1.

### v2

@@ VI.1 Bath parameters
In our numerical simulations, we first create a very large diamond lattice with the NV center in the center of this lattice and distribute P1 centers randomly on this lattice. P1-center spins sufficiently far away from the NV spin should have a negligible effect on the spin-coherence decay of the NV spin. Therefore, we limit the number of P1-center spins included in the simulation. We do this by introducing a bath radius $r_{\text{b}}$ which defines a sphere around the NV spin. Only spins within this sphere are considered dynamical.
@@
Figure S5: Evidence for the convergence of parameters for the pCCE(2,4) method applied to a large bath of P1-center electron spins in a quasi-two-dimensional configuration at a concentration of 1 ppm: a) the bath-radius $r_{\text{b}}$ and b) the dipole radius $r_{\text{d}}$. The values of $r_{\text{b}}$ and $r_{\text{d}}$ depend both on the layer thickness $L$ and the bath-spin concentration $\rho_{\text{P1}}$. In our calculations, we choose the values for the parameters $r_{\text{b}}$ and $r_{\text{d}}$ corresponding to converged spin-coherence decay results, i.e., the spin-coherence decay does not change appreciably when increasing any of the parameters. For further details, see text.
The values of the above parameters $r_{\text{b}}$ and $r_{\text{d}}$ were determined in convergence studies. An example is illustrated in Fig. S5, where the spin-coherence decay $\langle M_{x}(2\tau)\rangle$ of the NV spin is shown for a P1-center spin bath at concentration 1 ppm using pCCE(2,4). The spin-coherence decay $\langle M_{x}(2\tau)\rangle$ is averaged over 100 randomly distributed spin systems and 20 mean-field spin configurations in internal averaging. The bath radius is increased in steps of 10 nm starting from 60 nm, see Fig. S5 a. The dipole radius is increased in steps of 10% starting from the initial value of 64.9 nm, see Fig. S5 b. In both calculations, the other parameter was kept constant at its initial value. The resulting graphs do not show any significant changes when increasing the parameter values. We thus use the values $r_{\text{b}}=60$ nm and $r_{\text{d}}=64.9$ nm in our simulations for this particular system. Within the radius $r_{\text{b}}=60$ nm, there are 140 bath spins, which corresponds to the number of P1 centers. Small fluctuations of results shown in Fig. S5 arise presumably due to a finite number of samples for the mean-field average. The chosen parameter values are similar to those used in other studies of NV spins in P1-center baths for CCE2, see for example Ref. [50].
When considering other concentrations $\rho_{\text{P1}}$, the bath radius $r_{\text{b}}$ is chosen such that there are at least 140 spins, i.e., P1 centers, within the radius. To assure partitions of the same size within each P1-bath subgroup (discussed above), further spins are added to each subgroup until the total number of spins is divisible by the partition size $K$.
@@
where the last term on right-hand-side makes $r_{\text{d}}$ larger for quasi-two-dimensional layers. We also use the same values of $r_{\text{b}}$ and $r_{\text{d}}$ when leaving out the hyperfine interaction in the P1 centers.
Figure S6: Convergence of the pCCE method with respect to the partition size $K$ and the number of samples for the internal mean-field average for different spin baths. a) An electron spin-1/2 bath in a quasi-two-dimensional layer of thickness $L=5$ nm at a concentration of 1 ppm. At larger $K$, results for different numbers of samples for the internal average are also shown to demonstrate the convergence with the number of samples. b) A P1-center spin bath in a quasi-two-dimensional layer ($L=30$ nm) at a concentration of $\rho_{\text{P1}}=1$ ppm. Fluctuation of results of pCCE(2,3) likely arise due to insufficient internal mean-field averaging. In a and b, the results for $K=2$ and $K=3$ were averaged internally over 100 and 50 mean-field samples, respectively.
@@ VI.2 Partition size $K$ and the number of samples for the mean-field average
We first consider a quasi-two-dimensional system of electron spins-1/2 with layer thickness $L=5$ nm and concentration of 1 ppm and calculate the spin-coherence decay $\langle M_{x}(2\tau)\rangle$. Figure S6 a shows results for different partition sizes $K$ and different number of samples for the internal mean-field average. The results indicate that the convergence is achieved for partition size $K=4$ and 20 samples for the internal mean-field average.
@@
The decay of $\langle M_{x}(2\tau)\rangle$ becomes faster for larger $K$. This implies that, when using CCE2, the resulting $T_{2}$ time is too large in the quasi-two-dimensional case. On the other hand, when considering three-dimensional baths of electron spins-1/2, the pCCE(2,$K$) results for all $K\geq 1$ demonstrate similar behavior (not shown here) which indicates that already pCCE(2,1) (equivalent to CCE2) provides accurate results. For the sake of consistency, we use pCCE(2,4) also for these systems with the same number of samples for the internal average.
@@
For the P1-center spin baths, no significant difference between pCCE(2,1) and pCCE(2,4) was observed for all layers, as indicated in Fig. S6 b. We attribute this to the strong suppression of flip-flop transitions in the bath due to the hyperfine interaction, see the first section of this supplemental material. The CCE2 method thus yields sufficiently accurate results for P1-center spin baths. However, this is strictly speaking valid only for the spin-coherence decay averaged over many different random spin distributions. This does not imply that, for each individual spin system, CCE2 yields sufficiently accurate results, cf. Ref. [55]. Using pCCE, accurate results can also be obtained for individual systems. It is also noteworthy, that, for establishing the validity of CCE2 approach for P1-center spin baths, the extensive use of the pCCE method was necessary.
VII Fitting the stretched-exponential function
@@
Similar behavior is observed for the baths of P1 centers without the hyperfine interaction, i.e., baths of electron spins-1/2. The resulting curves and corresponding fits are shown in Fig. S8. The transitions from three-dimensional to quasi-two-dimensional systems occur at lower layer thickness compared to the P1-center spin bath. The reason for this is the absence of the suppression of the flip-flop transitions by the hyperfine interaction and, hence, different spin dynamics in the bath. Therefore, in Table 1 of the main text, we considered for the quasi-two-dimensional configuration for P1 centers a layer thickness of $L=30\,$nm at 2 ppm and 1 ppm and $L=60\,$nm at 0.1 ppm, and, for P1 centers without the hyperfine interaction, $L=5\,$nm at 2 ppm and $L=10\,$nm at 0.1 ppm and 1 ppm.
Figure S8: Illustration of the fitting of the stretched-exponential function in Eq. (2) to the numerical results obtained for P1 centers without the hyperfine interaction for three-dimensional (a) and quasi-two-dimensional (b) diamond layers, cf. Table 1 of the main article. The layer thickness $L$ and the P1-center concentration $\rho_{\text{P1}}$ are indicated in the legends. The fitting parameters are the stretched-exponential parameter $p$ and the $T_{2}$ time. The time window for the fit was adjusted for each individual curve separately.
@@
The above three regimes complicate the choice for a suitable region for the fit of the short-time decay of $\langle M_{x}(2\tau)\rangle$. Thus, it is difficult to reliably identify the beginning and the end of this region such that the obtained slope of the curves corresponding to the stretched-exponential parameters $p$ together with the characteristic $T_{2}$ time must be considered an approximation.
VIII Estimation of errors
The P1-center spin baths considered here are sparse, which means that the P1-center concentration $\rho_{\text{P1}}$ is much smaller than the carbon-atom concentration in pure diamond. This leads to large variations of the spatial P1-center spin distribution and, hence, to large variations in the spin-coherence decay. In Figure S9, we show all 100 individual spin-coherence decays together with the mean value (black dots) for a quasi two- (b) and a three-dimensional (a) spin bath. The error bars indicate the standard deviation at the individual times. The standard-deviation of the mean is by the factor $\sqrt{100}=10$ smaller and, hence, comparable or smaller than the point size used in the figure.
Figure S9: Spin-coherence decay for each individual spatial bath-spin distribution is shown for (a) a three-dimensional and (b) a quasi-two-dimensional P1-center spin bath. The mean is indicated by the black dots and the error bars indicate the standard deviation resulting from the large variation of the decay curves.
To estimate the errors of the $T_{2}$ time and the stretched-exponential parameter $p$, we consider the covariance matrix obtained from the linear fit to the spin-coherence decay in a double-logarithmic plot as detailed in the previous section. From this covariance matrix, the standard deviation of the fitting parameters is calculated. In this fit, the standard deviation of the mean values of each data point mentioned above is included. Further, we add to the error 5 % of the fitted values to account for the error arising from the choice of the time window for the fitting. The resulting error bars are shown in Fig. 4 of the main article.